OpenStudy (anonymous):

A rectangle has an area of 12 square centimeters and a perimeter of 16 centimeters. Which of the following could be its dimensions? 2 cm and 6 cm 3 cm and 4 cm 1.5 cm and 8 cm 1 cm and 12 cm

OpenStudy (anonymous):

i think its 2 and 6

OpenStudy (anonymous):

@imqwerty

OpenStudy (thehumantorch10):

the correct dimensions will solve both of these equations simultaneously, bh=12 , 2(b+h)=16 ie the dimensions will multiply to give twelve, and add to give 8

OpenStudy (ag23):

it is 2 and 6

OpenStudy (thehumantorch10):

you are right

OpenStudy (thehumantorch10):

👍

imqwerty (imqwerty):

let the lenth and breath of rectangle be l and b l x b = 12 l=12/b equation 1 and 2(l+b)=16 l+b=8 l=8-b equation2 eq1 = eq 2 12/b = 8-b 12=8b-b^2 b^2 -8b +12 = 0 (b-6)(b-2)=0 if b=6 l=2 if b=2 l=6

OpenStudy (anonymous):

<3